### v1

# Improving Generalization of Universal Adversarial Perturbation via Dynamic Maximin Optimization

Abstract
Deep neural networks (DNNs) are susceptible to universal adversarial perturbations (UAPs). These perturbations are meticulously designed to fool the target model universally across all sample classes. Unlike instance-specific adversarial examples (AEs), generating UAPs is more complex because they must be generalized across a wide range of data samples and models. Our research reveals that existing universal attack methods, which optimize UAPs using DNNs with static model parameter snapshots, do not fully leverage the potential of DNNs to generate more effective UAPs. Rather than optimizing UAPs against static DNN models with a fixed training set, we suggest using dynamic model-data pairs to generate UAPs. In particular, we introduce a dynamic maximin optimization strategy, aiming to optimize the UAP across a variety of optimal model-data pairs. We term this approach DM-UAP. DM-UAP utilizes an iterative max-min-min optimization framework that refines the model-data pairs, coupled with a curriculum UAP learning algorithm to examine the combined space of model parameters and data thoroughly. Comprehensive experiments on the ImageNet dataset demonstrate that the proposed DM-UAP markedly enhances both cross-sample universality and cross-model transferability of UAPs. Using only 500 samples for UAP generation, DM-UAP outperforms the state-of-the-art approach with an average increase in fooling ratio of 12.108%.
Codehttps://github.com/yechao-zhang/DM-UAP
1 Introduction
Deep Neural Networks (DNNs) represent the cutting edge of computer vision (Krizhevsky, Sutskever, and Hinton 2012; Simonyan and Zisserman 2014; He et al. 2016). A key strength of DNNs is their ability to identify subtle patterns beyond human perception. However, this capability of DNNs can unintentionally enable adversarial attacks, where deceptively minor adversarial perturbations of the input—imperceptible to humans—create adversarial examples (AE) that lead to mispredictions of DNNs (Szegedy et al. 2014).
Adding to this concern is the emergence of a notable class of adversarial techniques known as universal adversarial perturbation (UAP), which are designed to induce mispredictions against universal input samples (Moosavi-Dezfooli et al. 2017; Khrulkov and Oseledets 2018; Poursaeed et al. 2018; Shafahi et al. 2020). In addition to this characteristic of cross-sample universality, UAPs could also possess another element of generalization in adversarial attacks, termed cross-model transferability. This denotes the capacity of (AEs) to preserve their efficacy across different DNN models (Hu et al. 2022; Li et al. 2024). Together, these capabilities enable UAPs to facilitate the generation of transferable AEs at scale, compromising various DNN architectures and multiple machine learning tasks simultaneously (Zhong and Deng 2022; Xie et al. 2021; Ding et al. 2021; Wang et al. 2023; Zhou et al. 2023a, b, 2024). Consequently, UAPs have garnered significant attention and have become a focal point of extensive research in the field.
Recent research aimed at improving the generalization of UAPs has explored different aspects. Some focus on analyzing the training data used for UAP generation (Zhang et al. 2020b; Li et al. 2022), while others reinterpret the gradient aggregation within and between the training batch (Shafahi et al. 2020; Liu et al. 2023). In addition, refinements to training loss have also been explored (Zhang et al. 2021). However, these existing works fall short by only refining UAPs against fixed DNN parameters, thereby limiting the generalization to static instances of the model without fully harnessing the adaptive capabilities inherent in the model.
In this work, we highlight that utilizing the ever-evolving optimized model parameter configurations can enable the generation of stronger and more transferable UAPs. The intuition behind this is that exposing the generation process to a wider range of model variance, even within a fixed architecture, fosters UAPs to capture a wider spectrum of vulnerabilities between models. This, in turn, allows UAPs to generalize to models on which they were not directly optimized. Furthermore, we embed the objectives of cross-sample universality and cross-model transferability into a unified maximin formulation, which involves the composite inner minimization of the model parameters and training data and an outer maximization of the UAP. In other words, we generate UAPs on a wide range of parameter-data pairs that have jointly minimized the classification loss w.r.t. their respective labels. In this way, the generated UAP is optimized to maximize the loss in more antagonistic scenarios from a combined landscape of both parameter and input space.
To effectively address the optimization challenges associated with this maxmini formulation, we develop an iterative max-min-min optimization framework with dynamic adaptation designs. We decouple the inner composite optimization into a two-stage min-min optimization process, where we dynamically optimize a parameter-data pair within each minibatch iteration. Additionally, we introduce curriculum UAP learning, in which we first proceed the maximization against “easy” parameter-data pairs, and gradually increase their optimization spaces along with the training epoch, to ensure a smoother UAP optimization. In general, the main contributions of our work can be summarized as follows:
  • We propose a novel maximin formulation, which adapts the model parameters during the optimization process to improve the generalization of UAP. To the best of our knowledge, this is the first exploration of the model parameter landscape for UAP generation.
  • To effectively solve the maximin formulation, we propose a dynamic iterative max-min-min optimization framework, which decouples the composite minimization into a two-stage min-min optimization and leverages curriculum learning to ensure a smooth UAP update.
  • Extensive experiments on the ImageNet dataset demonstrate the superior generalization of UAPs generated by the proposed DM-UAP compared to the state-of-the-art methods under various attack settings.
2 Related Work
2.1 Instance-specific Attacks
Instance-specific attacks generate AEs by adding perturbations to individual samples. In a white-box setting, the adversary has full knowledge of the target model (architecture and parameters), constructing AEs using directly computed gradients, in a single-step (Goodfellow, Shlens, and Szegedy 2014) or iterative manner (Madry et al. 2017). In a black-box setting, the adversary lacks direct access to the target model. Alternatively, AEs are crafted against surrogate models, expecting effectiveness against the black-box target model. Since the discovery of instance-specific AE (Szegedy et al. 2013), research has extensively studied their transferability, exploring loss function refinement (Zhao, Liu, and Larson 2021), input transformations (Xie et al. 2019), and model ensembles (Zhang et al. 2024) to improve the effectiveness.
2.2 Universal Adversarial Attacks
The seminal work (Moosavi-Dezfooli et al. 2017) of UAPs, denoted as UAP, proposed aggregating perturbation vectors obtained from instance-specific attack methods to generate UAPs. Subsequent work, such as GAP (Poursaeed et al. 2018) and NAG (Mopuri et al. 2018), introduced the use of generative models for UAP generation.
In a later development, SPGD (Shafahi et al. 2020) formulated the generation of UAPs as a maximization of the average prediction loss on universally perturbed training data and used the stochastic gradient method with mini-batch training to solve it. Since then, the mini-batch training paradigm has become a common approach to UAP generation in subsequent works (Co et al. 2021; Zhang et al. 2020a; Benz et al. 2020; Liu et al. 2023; Hu et al. 2021). Among these, DF-UAP (Zhang et al. 2020b) found that UAPs dominate DNN prediction, and the original samples behave somewhat like random noise after superimposing UAPs. This finding inspired AT-UAP (Li et al. 2022), a more robust UAP achieved by integrating image-specific attacks and universal attacks, to enhance such a dominant effect. Recently, SGA (Liu et al. 2023) proposed to aggregate the noisy gradients obtained from multiple small-batch as a gradient estimation of a large-batch. However, these methods have not yet sought to improve the generalization of UAPs from the perspective of manipulating the underlying model.
3 Methodology
In this section, we first introduce existing UAP problem formulations and present our novel maximin formulation. After that, we present the details of our proposed max-min-min optimization framework for solving this formulation.
3.1 Preliminaries
Suppose $\mathbf{S}$ contains a set of samples $(x,y)\in\mathcal{D}$ drawn from the data distribution $\mathcal{D}$, where $x\in\mathcal{X}$ represents the image feature and $y\in\mathcal{Y}$ denotes its corresponding label. The DNN model is depicted as $f_{\theta}$, where $f$ is a DNN function and $\theta$ represents the model parameters. In general, $\theta$ are optimized through empirical risk minimization (ERM) as follows:
$\min_{\theta}\frac{1}{\|\mathbf{S}\|}\sum_{i=1}^{\|\mathbf{S}\|}\mathcal{L}\left(f_{\theta}\left(x_{i}\right),y_{i}\right),$ (1)
where $\mathcal{L}$ is the loss function (e.g., cross-entropy) for training.
Empirical Risk Maximization. In a universal adversarial attack, the objective is to craft a single perturbation $\delta$ to fool the well-trained DNN model $f_{\theta}$ for most images within $\mathcal{X}$. In practice, existing work aims to optimize $\delta$ which maximizes the averaged prediction loss for universally perturbed data as follows:
$\max_{\|\delta\|_{\infty}\leq\epsilon}\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}\left(f_{\theta}\left(x_{i}+\delta\right),y_{i}\right),$ (2)
where $x_{i},...,x_{n}\in\mathcal{X}$ are training data used for optimizing the $\delta$, $y_{i}$ is the corresponding label of the unperturbed input $x_{i}$ predicted by $y_{i}=\arg\max f_{\theta}\left(x_{i}\right)$, and UAP $\delta$ is a small perturbation constrained by $\ell_{\infty}$-norm of $\epsilon$.
Figure 1: The illustration of different optimization flows: a) Eq. 2 use the original data and model for UAP generation; b) Eq. 3 use the model and original inputs to obtain optimized inputs, then use model and optimized inputs for UAP generation; c) Eq. 5 use original model and original inputs to obtain optimized models and inputs, then use them for UAP generation.
Optimal-Data UAP Maximin. In addition to the standard empirical risk maximization, (Li et al. 2022) introduces a maximin optimization that leverages image-specific adversarial attacks to produce more effective training inputs. This leads to crafting a UAP based on a tailored training dataset. Formally, this maximin optimization is depicted as follows:
$\begin{gathered}\max_{\|\delta\|_{\infty}\leq\epsilon}\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}(f_{\theta}(x_{i}^{*}+\delta),y_{i}),\\ \text{s.t. }\forall i,x_{i}^{*}=\arg\min_{\|x^{\prime}-x_{i}\|_{2}\leq r}\mathcal{L}\left(f_{\theta}(x^{\prime}),y_{i}\right).\end{gathered}$ (3)
This maximin optimization searches for optimal data point $x_{i}^{*}$ in the $r$-bounded $\ell_{2}$-norm vicinity of each original input $x_{i}$, which minimizes the loss for its corresponding label $y_{i}$ first. Then the UAP $\delta$ is optimized to maximize the loss over these optimal data points. Compared to Eq. 2, this formulation constructs antagonistic training data for maximization, compelling the UAP to avoid trivial solutions that may not effectively deceive the model $f_{\theta}$.
3.2 Problem Formulation
In contrast to previous formulations (Eqs. 2 and 3) that optimize UAP against a static model $f_{\theta}$—whether for a direct white-box attack or through a surrogate model in transfer-based attacks—we propose optimizing across multiple variants of the model parameter $\theta$ to enhance the potency of the UAP.
Optimal-Parameters UAP Maximin. Inspired by the optimal-data UAP maximin concept in Eq. 3, we introduce a parallel optimization as follows:
$\begin{gathered}\max_{\|\delta\|_{\infty}\leq\epsilon}\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}(f_{\theta_{i}^{*}}(x_{i}+\delta),y_{i}),\\ \text{s.t. }\forall i,\theta_{i}^{*}=\arg\min_{\|\theta^{\prime}-\theta\|_{2}\leq\rho}\mathcal{L}\left(f_{\theta^{\prime}}\left(x_{i}\right),y_{i}\right).\end{gathered}$ (4)
Unlike Eq. 3, this optimization finds the optimal model parameters $\theta_{i}^{*}$ that minimize the loss for each data point $(x_{i},y_{i})$ within a $\rho$-bounded $\ell_{2}$-norm vicinity of the original parameters $\theta$. The UAP $\delta$ is then optimized to maximize the average loss across all perturbed data points $(x_{i}+\delta,y_{i})$, evaluated using their respective model variants $f_{\theta_{i}^{*}}$. This ensures UAP effectively increases loss across different model variations, promoting a robust adversarial effect expected to generalize better across models.
Optimal-Composition UAP Maximin. Building on the optimal-parameters UAP maximin, we propose a more advanced formulation that includes a composite inner minimization over both model parameters and data points:
$\begin{gathered}\max_{\|\delta\|_{\infty}\leq\epsilon}\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}(f_{\theta_{i}^{*}}(x_{i}^{*}+\delta),y_{i}),\\ \text{s.t. }\forall i,(\theta_{i}^{*},x_{i}^{*})=\arg\min_{\begin{subarray}{c}\|\theta^{\prime}-\theta\|_{2}\leq\rho\\ \|x^{\prime}-x_{i}\|_{2}\leq r\end{subarray}}\mathcal{L}(f_{\theta^{\prime}}(x^{\prime}),y_{i}).\end{gathered}$ (5)
In this context, each $(\theta_{i}^{*},x_{i}^{*})$ pair is simultaneously optimized to reduce the loss for $y_{i}$, while remaining within their respective neighborhoods defined by $\rho$ and $r$. This sets up a more challenging condition for the UAP $\delta$, which requires it to maximize loss against the best-modified models and their corresponding inputs. Both Eqs. 3 and 4 represent special cases of this formulation. Fig. 1 illustrate the optimization flow of Eqs. 2, 3 and 5.
3.3 Dynamic Maximin UAP: An Iterative Max-Min-Min Optimization Framework
Tackling Eq. 5 is a non-trivial task. First, the minimization process involves a dual-layered optimization that requires accurate navigation through two high-dimensional spaces. Second, the maximization phase must be flexible enough to cope with a wide range of combined landscapes concerning both data and the model parameters, which poses greater optimization challenges compared to Eqs. 2, 3 and 4. To overcome these difficulties, we introduce an iterative max-min-min optimization framework. This framework separates the composite minimization into a two-stage min-min optimization and employs curriculum learning for a smoother UAP maximization, as detailed below.
Dynamic Strategy for mini-batch Training. Intuitively, Eq. 5 suggests a straightforward implementation: find all optimal pairs $(x_{i}^{*},f_{\theta_{i}^{*}})$, use them as a fixed training set to update UAP through mini-batch training with stochastic gradient ascent. However, this is highly inefficient due to the significant storage and memory required for forward-backward propagation across multiple models $f_{\theta_{i}^{*}}$. To address this, we adopt a dynamic strategy: during each mini-batch iteration, we optimize a mini-batch of images with a shared model on the fly, i.e., each random mini-batch of images $X_{B}$ with the initial model $f_{\theta}$ induces a composition $(X_{B}^{*},f_{\theta^{*}})$. After that, the perturbed mini-batch data $X^{*}_{B}+\delta$ are propagated forward and backward in parallel through $f_{\theta^{*}}$ to obtain the update gradient for $\delta$.
Sequential Model and Data Optimization. As discussed earlier, within each mini-batch iteration, we decouple the composite minimization of the data and model parameters into sequential min-min optimization steps. Specifically,
  • We first perform model optimization on the mini-batch images $X_{B}$ to obtain the optimized model $f_{\theta^{*}}$.
  • Then, we use the optimized model $\theta^{*}$ to perform data optimization to obtain $X_{B}^{*}$.
The rationale for this “optimize model first, data second” approach is to ensure that the model is always updated using unperturbed data. This is because adversarially constructed images $X_{B}^{*}$ may not belong to the correct feature distribution $\mathcal{X}$. Overfitting the model to adversarially constructed images $X_{B}^{*}$ could result in significant underfitting of the original distribution (Dong et al. 2020; Xu et al. 2022).
Model Optimization. To ensure the effectiveness of resulting UAP against the original model $f_{\theta}$, the optimized model parameters $\theta^{*}$ should not deviate much from the origin $\theta$. However, standard gradient optimizers cannot guarantee precise adherence to the $\ell_{2}$-norm constraint. To mitigate this, we leverage the normalized gradient of parameters when updating $\theta^{*}$. Concretely, we update $\theta^{*}$ as follows:
$\theta^{*}\leftarrow\theta^{*}-\alpha_{m}\frac{\nabla_{\theta}\frac{1}{B}\mathcal{L}(f_{\theta^{*}}(X_{B}),Y_{B})}{\|\nabla_{\theta}\frac{1}{B}\mathcal{L}(f_{\theta^{*}}(X_{B}),Y_{B})\|_{2}},$ (6)
where $Y_{B}$ are the corresponding labels of images $X_{B}$, $\alpha_{m}$ is the step size of model optimization. Given the adversarial budget $\rho_{t}$ at the $t$-th epoch, we update $\theta^{*}$ for $K_{m}$ steps and set the step size $\alpha_{m}$ as $\frac{\rho_{t}}{K_{m}}$ to ensure that $\theta^{*}$ remain within the $\rho_{t}$-bounded neighborhood of $\theta$.
Data Optimization. As discussed earlier, the data optimization is conducted in condition to the optimized model $f_{\theta^{*}}$ inside each mini-batch iteration. Specifically, we update the data using the standard $\ell_{2}$-norm targeted PGD as follows:
$x^{*}\leftarrow\Pi_{r_{t}}(x^{*}-\alpha_{d}\cdot\frac{\nabla_{x}\mathcal{L}(f_{\theta^{*}}(x^{*}),y)}{\|\nabla_{x}\mathcal{L}(f_{\theta^{*}}(x^{*}),y)\|_{2}}),$ (7)
where $x^{*}$ denotes a sample in mini-batch $X_{B}^{*}$, and $\Pi(\cdot)$ is the projection function that ensures the data perturbation does not exceed the budget of $r_{t}$. We set the step size of data optimization $\alpha_{d}$ as $1.25\times\frac{r_{t}}{K_{d}}$. Eq. 7 is executed for all the samples in $X_{B}^{*}$ in parallel, and repeated for $K_{d}$ steps.
Input: Data $X={x_{1},\ldots,x_{n}}$, model $f$ with parameters $\theta$, number of epochs $T$, mini-batch size $B$.
Input: UAP maximum perturbation magnitude $\epsilon$, initial learning rate $\gamma$.
Input: Model maximum neighborhood size $\rho$, number of model optimization steps $K_{m}$.
Input: Data maximum neighborhood size $r$, number of data optimization steps $K_{d}$.
Output: Universal Adversarial Perturbation (UAP) $\delta$
1
2Initialize $\delta\sim\mathcal{U}(-\epsilon,\epsilon)$;
3for $t=1$ to $T$ do
4       Compute the perturbation magnitude $\rho_{t}=t\times\frac{\rho}{T}$ and step size $\alpha_{m}=\frac{\rho_{t}}{K_{m}}$ for model optimization in epoch $t$;
5      Compute the perturbation magnitude $r_{t}=t\times\frac{r}{T}$ and step size $\alpha_{d}=1.25\times\frac{r_{t}}{K_{d}}$ for data optimization in epoch $t$;
6      for mini-batch $X_{B}\in X$ do
7
8            $Y_{B}=\arg\max f_{\theta}(X_{B})$;
9
10            Initialize $\theta^{*}\leftarrow\theta$;
11             Initialize $X^{*}_{B}\leftarrow X_{B}$ ;
12
13            for $k=1$ to $K_{m}$ do
14
15                  Update model parameters with Eq. 6;
16
17             end for
18
19            for $k=1$ to $K_{d}$ do
20
21                  Update data with Eq. 7 (in parallel);
22
23             end for
24
25            Update UAP with Eq. 8 ;
26
27       end for
28
29 end for
return $\delta$;
Algorithm 1 Dynamic Maximin UAP (DM-UAP) with Curriculum Learning
Curriculum UAP Learning. The outer maximization of updating UAP is challenging as it requires accumulating gradients from various parameters and inputs, all of which are constantly changing in high-dimensional neighborhoods. To address this, we leverage the concept of curriculum learning (Bengio et al. 2009). We adopt a progressive approach by initiating inner-min-min optimization of $\theta^{*}$ and $X_{B}^{*}$ in smaller neighborhoods during the early training epochs. Then, we gradually increase the neighborhood sizes to encompass larger regions over the course of the training. In each epoch $t$ out of a total of $T$ epochs, we set the neighborhood sizes as $\rho_{t}=t\times\frac{\rho}{T}$ and $r_{t}=t\times\frac{r}{T}$, respectively.
To better update UAP in a dynamic environment, instead of using SGD with the sign function as in previous works (Liu et al. 2023), we use Adam as the optimizer. Since Adam adjusts learning rates and momentum to adapt to the optimization landscapes, which makes it suitable for such non-stationary objectives, the UAPs are updated adaptively to the evolving dynamics of the model and data. Formally, the UAP is updated in each iteration as follows:
$\begin{gathered}\delta=Adam(\nabla_{\delta}\mathcal{L}(f_{\theta^{*}}(X_{B}^{*}+\delta),Y_{B}),\gamma)\\ \delta=\min(\max(\delta,-\epsilon),\epsilon)\end{gathered}$ (8)
where $\gamma$ is the initial learning rate, and the min-max operation is conducted to ensure the valid $\ell_{\infty}$-norm constraint on the generated UAP. The entire framework of our proposed DM-UAP method is summarized in Algorithm 1.
4 Experiment
Setup: Following (Moosavi-Dezfooli et al. 2017; Liu et al. 2023), we randomly select 10 images from each category in the ImageNet training set, resulting in a total of 10,000 images, for UAP generations. In addition, we also consider a data-limit setting, in which only 500 random images from the training set are sampled. Aligning with previous work, we evaluate our method on the ImageNet validation set, which contains 50,000 images, using classical pre-trained CNN models AlexNet, GoogleNet, VGG16, VGG19, and ResNet152 as target models.
Evaluation Metrics: We employ the widely used fooling ratio metric, which calculates the variation proportion of model predictions when applying the UAP, for evaluation.
Comparative Methods: In the white-box attack scenario, we compare our method with the following existing approaches: UAP, NAG, GAP, DF-UAP, Cos-UAP, AT-UAP, and SGA. For other settings, we compare DM-UAP with UAP, GAP, SPGD, AT-UAP, and SGA. SPGD is considered the baseline, as AT-UAP, SGA, and our proposed DM-UAP all build upon its standard mini-batch training. Note that SGA is the current state-of-the-art method.
Hyper-parameters Setting: We set the maximum perturbation budget $\epsilon$ of all methods as $10/255$. Following SGA (Liu et al. 2023), the number of training epochs $T$ is 20, and the batch size $B$ is 125. The step numbers for inner model optimization $K_{m}$ and data optimization $K_{d}$ in our method are both 10, with default neighborhood size $\rho=1$ and $r=32$.
Method ANet GNet V16 V19 R152 Avg
UAP 93.30 78.90 78.30 77.80 84.00 82.46
NAG 96.44 90.37 77.57 83.78 87.24 87.08
GAP - 82.70 83.70 80.10 - 82.17
DF-UAP 96.17 88.94 94.30 94.98 90.08 92.89
Cos-UAP 96.50 90.50 97.40 96.40 90.20 94.20
AT-UAP 97.01 90.82 97.51 97.56 91.52 94.88
SGA 96.99 90.64 97.83 96.56 92.86 94.98
Ours 97.19 93.28 98.43 97.81 92.90 95.92
Table 1: The fooling ratio (%) in the white-box setting by various UAP attack methods. The UAPs are crafted on the AlexNet, GoogleNet, VGG16, VGG19, and ResNet152.
4.1 Generalization Performance of UAPs
We perform universal adversarial attacks under the white-box and black-box settings respectively and evaluate the overall performance of our proposed DM-UAP with baselines on the ImageNet validation set.
White-box Attack. We present the results of white-box attacks on five models using our DM-UAP approach, compared with other methods in Tab. 1. For SGA, we used its official source code and followed the settings from (Liu et al. 2023) with 10,000 samples. For other methods, we used results from their respective papers. Our method achieves the highest attack performance across all models. DM-UAP shows a notable improvement of over 2% on GoogleNet. These indicate UAPs generated by our method can better generalize to unknown samples.
Black-box Attack. We evaluate transfer attacks with 10,000 and 500 training images. UAPs are generated for five considered models, and AEs are transferred between them. As shown in Tab. 3, DM-UAP outperforms others across all models. Compared to AT-UAP and SGA, DM-UAP consistently achieves the highest average fooling ratio improvement over SPGD for all surrogate models, while AT-UAP and SGA are sometimes inferior to SPGD. With 10,000 images, DM-UAP’s average fooling ratio improvement over SPGD ranges from 2.76% to 11.15%. With 500 images, DM-UAP’s improvement ranges from 8.22% to 45.19%, compared to AT-UAP’s -1.51% to 38.00% and SGA’s 0.77% to 22.55%. DM-UAP outperforms SGA with an average increase in fooling ratio of 12.108%.This manifests the importance of a dynamic model landscape proposed in our formulation in the limited samples scenario.
Method Samples ANet V16 GNet V19 R50 R152 Avg
SPGD 500 52.34* 58.45* 22.80 45.78 24.22 19.72 37.22
M-SPGD 72.38* 86.13* 33.14 70.46 35.78 28.12 54.34${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+17.12)}}}$
AT-UAP 80.41* 92.11* 42.28 80.29 43.59 33.49 62.03${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+24.81)}}}$
SGA 82.34* 90.70* 51.07 78.79 50.95 40.76 65.77${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+28.55)}}}$
M-SGA 85.93* 91.42* 47.69 78.41 48.77 37.95 65.03${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+27.81)}}}$
Ours 91.39* 92.83* 56.23 82.74 55.01 42.62 70.14${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+32.92)}}}$
SPGD 10000 71.02* 96.48* 50.21 89.48 53.90 42.75 67.31
M-SPGD 73.35* 97.47* 52.18 91.42 56.19 44.07 69.11${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+1.80)}}}$
AT-UAP 82.33* 97.23* 56.83 91.22 58.58 46.81 72.17${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+4.86)}}}$
SGA 82.38* 97.30* 60.46 91.49 62.21 51.26 74.18${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+6.87)}}}$
M-SGA 80.48* 97.83* 60.23 92.46 61.86 50.51 73.89${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+6.58)}}}$
Ours 91.35* 96.91* 66.34 91.33 63.88 51.48 76.88${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+9.57)}}}$
Table 2: The fooling ratio (%) in the ensemble-model setting by different UAP generation methods. The UAPs are crafted on the ensemble models, i.e., AlexNet and VGG16.
4.2 Scalability Performance of UAPs
In this subsection, we analyze the scalability performance of the proposed method from various aspects, including ensemble model setting, diverse sample setting, Transformer-to-CNN setting, and the attack-under-defense setting.
Model Method 10,000 samples 500 samples
AlexNet GoogleNet VGG16 VGG19 ResNet152 Average AlexNet GoogleNet VGG16 VGG19 ResNet152 Average
AlexNet UAP 84.28* 31.12 39.17 37.24 22.11 42.78 48.67* 18.13 23.07 22.59 14.28 25.35
GAP 88.55* 35.52 52.85 49.22 28.81 50.99 83.96* 32.99 48.26 45.17 26.98 47.47
SPGD 96.30* 54.16 61.39 58.89 36.78 61.50 89.20* 40.79 51.15 49.53 28.05 51.74
AT-UAP 96.74* 48.86 62.23 58.80 33.36 60.00(-${}_{\textrm{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-1.50)}}}$ 93.80* 31.78 51.98 48.57 25.00 50.23(-${}_{\textrm{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-1.51)}}}$
SGA 96.99* 46.62 65.57 59.81 34.30 60.66(-${}_{\textrm{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-0.84)}}}$ 94.91* 34.88 55.27 50.85 26.64 52.51${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+0.77)}}}$
Ours 97.19* 53.95 68.20 63.03 38.91 64.26${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+2.76)}}}$ 96.13* 48.11 63.28 58.38 33.92 59.96${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+8.22)}}}$
GoogleNet UAP 39.55 55.01* 49.82 49.11 29.66 44.63 23.83 15.91* 19.45 18.94 11.78 17.98
GAP 53.31 80.21* 72.56 70.62 49.85 65.31 37.59 33.32* 35.63 35.41 20.81 32.55
SPGD 50.47 86.09* 66.79 65.93 44.66 62.79 31.28 55.35* 29.50 28.69 17.89 32.54
AT-UAP 54.13 92.55* 79.13 76.54 53.51 71.17${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+8.38)}}}$ 49.49 80.63* 65.03 63.23 41.14 59.90${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+27.36)}}}$
SGA 59.22 88.46* 77.10 74.80 54.97 70.91${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+8.12)}}}$ 50.77 68.10* 60.18 59.13 37.28 55.09${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+22.55)}}}$
Ours 57.62 93.28* 81.45 80.05 57.29 73.94${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+11.15)}}}$ 57.29 88.07* 76.88 74.57 51.43 69.65${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+37.11)}}}$
VGG16 UAP 33.33 36.19 75.36* 64.09 31.33 48.06 22.87 14.97 26.67* 22.84 12.72 20.01
GAP 35.90 50.05 84.59* 76.49 38.97 57.20 34.65 26.39 57.82* 44.69 21.39 37.00
SPGD 40.67 43.46 92.81* 83.18 44.44 60.91 34.57 24.10 73.10* 56.04 21.06 41.77
AT-UAP 43.72 42.39 96.68* 88.13 34.97 61.18${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+0.27)}}}$ 43.34 33.48 90.13* 75.06 28.84 54.17${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+12.40)}}}$
SGA 44.48 52.53 97.83* 92.36 48.30 67.10${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+6.19)}}}$ 42.85 41.96 92.64* 80.83 36.24 58.90${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+17.13)}}}$
Ours 49.35 57.18 98.43* 94.12 49.97 69.81${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+8.90)}}}$ 45.99 49.96 96.78* 89.03 43.57 65.07${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+23.30)}}}$
VGG19 UAP 33.98 36.62 64.57 74.77* 30.48 48.08 23.59 15.26 25.13 26.09* 12.77 20.57
GAP 45.23 43.69 73.89 82.31* 30.02 55.03 38.86 30.36 52.12 60.83* 23.19 41.07
SPGD 40.76 47.84 84.26 92.90* 45.43 62.24 33.38 22.30 52.50 65.49* 19.79 38.69
AT-UAP 45.42 43.72 90.61 95.47* 40.05 63.05${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+0.81)}}}$ 42.82 34.20 77.99 88.02* 29.89 54.58${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+15.89)}}}$
SGA 45.89 55.53 92.62 96.56* 49.90 68.10${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+5.86)}}}$ 42.70 42.90 84.42 90.88* 36.69 59.52${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+20.83)}}}$
Ours 49.81 59.91 95.51 97.81* 55.75 71.76${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+9.52)}}}$ 47.24 50.97 91.33 96.55* 41.57 65.53${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+26.84)}}}$
ResNet152 UAP 35.15 36.04 49.19 47.50 57.36* 45.05 27.28 18.06 25.02 24.10 17.97* 22.49
GAP 47.70 56.08 68.63 66.51 73.80* 62.54 39.60 38.23 46.90 46.24 46.66* 43.53
SPGD 46.12 55.01 77.33 74.19 90.33* 68.60 30.23 19.01 27.33 26.40 18.87* 24.37
AT-UAP 47.69 57.94 77.88 75.09 92.07* 70.13${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+1.53)}}}$ 44.67 47.62 71.56 67.51 80.49* 62.37${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+38.00)}}}$
SGA 49.30 62.40 80.53 78.18 92.86* 72.65${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+4.05)}}}$ 40.53 35.57 49.45 47.51 43.01* 43.21${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+18.84)}}}$
Ours 50.09 63.50 81.49 79.10 92.90* 73.42${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+4.82)}}}$ 51.26 59.17 78.05 75.40 83.92* 69.56${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+45.19)}}}$
Table 3: The fooling ratio (%) on five models in the black-box setting by different UAP attack methods. The UAPs are crafted on AlexNet, GoogleNet, VGG16, VGG19, and ResNet152, respectively. We conduct the evaluation using 10,000 and 500 images for training, respectively. The respective last column represents the average improvement or deterioration.
Ensemble-Model Setting. Following (Liu et al. 2023), we implement the model ensemble method using the averaged loss functions of two models, i.e., AlexNet and VGG16. In this experiment, we still use SPGD as the baseline method, and compare the improvement of DM-UAP with those of others. For more comparison, we also integrate SPGD and SGA with the momentum (Dong et al. 2018) method, denoted as M-SPGD and M-SGA. In addition to the commonly used five models, we also test the transferability on ResNet50. The results are reported in Tab. 2. DM-UAP still outstrips the others by a clear superiority, which verifies the effectiveness of our dynamic maximin formulation still suffices in such a more complex optimization landscape.
Diverse-Sample Setting. Furthermore, we investigate the impact of varying the number of training samples on the attack performance. The results are depicted in Fig. 2.
Figure 2: Average fooling ratio (%) on five models in the diverse-sample training scenarios, using VGG19.
Our method consistently achieves the best performance with any number of training samples. However, we observe that once the number of samples exceeds a certain threshold, the corresponding increase in attack performance plateaus. The phenomenon underscores the limitations of excessive increase in the number of training samples.
Transformer-to-CNN Setting. Unlike CNN models, transformer models employ self-attention mechanisms rather than convolutional blocks. Studies indicate that transformer models demonstrate reduced cross-model transferability to other architectures, particularly CNN. We adapt transformer models from DeiT (Touvron et al. 2021) and ViT (Dosovitskiy et al. 2020) family as surrogate models and evaluate the cross-model performance of different UAP methods. We can see from Tab. 4, that compared to SPGD, AT-UAP and SGA, our method consistently achieves the highest fooling ratio against all black-box CNN models.
Model UAP Method ANet GNet V16 V19 R152 Avg
ViT-B SPGD 51.65 30.10 49.39 47.66 23.86 40.53
SGA 54.31 37.86 54.18 52.26 28.93 45.51${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+4.98)}}}$
AT-UAP 56.24 44.88 55.47 54.19 32.31 48.62${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+8.09)}}}$
Ours 59.01 51.63 62.35 61.62 38.31 54.58${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+14.05)}}}$
ViT-L SPGD 35.94 24.93 36.01 34.83 18.55 30.05
SGA 43.34 34.69 47.18 46.35 24.60 39.23${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+9.18)}}}$
AT-UAP 51.38 40.14 53.73 51.47 28.63 45.07${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+15.02)}}}$
Ours 54.94 46.55 59.50 56.26 32.52 49.95${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+19.90)}}}$
DeiT-S SPGD 34.38 21.76 41.09 37.47 19.53 30.85
SGA 48.35 40.69 49.63 47.43 28.37 42.89${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+12.04)}}}$
AT-UAP 54.29 46.59 57.36 55.11 32.21 49.11${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+18.26)}}}$
Ours 56.60 50.09 62.00 59.64 37.35 53.14${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+22.29)}}}$
DeiT-B SPGD 39.66 26.72 40.02 38.04 19.66 32.82
SGA 43.30 27.81 44.65 42.52 21.18 35.89${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+3.07)}}}$
AT-UAP 41.12 27.37 48.41 44.44 21.74 36.62${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+3.80)}}}$
Ours 46.32 29.56 51.26 47.12 23.22 39.50${}_{\textrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}(+6.68)}}}$
Table 4: Fooling ratio (%) in the Transformer-to-CNN setting using different generation methods.
(a) Curriculum Learning
(b) Optimization order
(c) Model neighborhood size
(d) Data neighborhood size
Figure 3: Ablation study on model and data optimization. “M” means “Model Optimization” and “D” means “Data Optimization”, “M-D” denotes “Model-Data Optimization”, “D-M” denote “Data-Model Optimization”. (a) Fooling ratios of DM-UAP with/without curriculum learning, i.e, increasing neighborhood sizes. (b) Fooling ratio by different optimization orders in white-box setting for five models. (c) Fooling ratios of DM-UAP with/without data optimization for different model neighborhood sizes. (d) Fooling ratios of DM-UAP with/without model optimization for different data neighborhood sizes.
4.3 Ablation Study
We conduct ablation studies to evaluate the impact of curriculum learning, the model and data optimization order, and optimization neighborhood sizes on the performance of our proposed framework, using 500 training images by default.
On the curriculum UAP learning. We examine the impact of curriculum learning by gradually increasing the model neighborhood size $\rho$ from 0 to 4 and the data neighborhood size $r$ from 0 to 32 throughout the training epochs. For comparison, fixed neighborhood sizes of $\rho=4$ and $r=32$ are also used. As shown in Fig. 3(a), the results underscore the importance of using an easy-to-hard antagonistic approach as in curriculum learning for optimizing UAP, which significantly enhances its effectiveness.
On the optimization order of model and data. We explore the influence of different optimization orders to highlight the “optimize model first, data second” sequencing adopted in DM-UAP. In Fig. 3(b), we use “Model-Data Optimization” to denote the process of 10 model optimization steps followed by 10 data optimization steps as in DM-UAP. In contrast, “Data-Model Optimization” reverses this order. “Alternating Optimization” means alternating single optimization steps between the model and data 20 times. “Without Optimization” means both model and data are not optimized, which reduces to the formulation of Eq. 2. The results show “Model-Data Optimization” consistently outperforms the “Without Optimization” baseline. “Data-Model Optimization” sometimes, while “Alternating Optimization” consistently underperforms. These confirm our hypothesis that the model should be optimized using unperturbed data.
On the model neighborhood size. We explore the influence of the maximum model neighborhood size, with data neighborhood size set as $r=32$ for “Model-Data Optimization”. We tested $\rho$ values of 1, 2, 4, 8, 10, 12, and 16, as depicted in Fig. 3(c). For comparison, the formulation of optimal-parameters UAP maximin in Eq. 4 is also assessed, denoted as “Model Optimization”. The results indicate that “Model-Data Optimization” consistently surpasses “Model Optimization”, and their optimal $\rho$ value differ.
On the data neighborhood size. Similarly, we explore the influence of the maximum data neighborhood size $r$, with the model neighborhood size set as $\rho=4$ for “Model-Data Optimization”. We test $r$ values of 1, 4, 8, 16, 32, 64, and 100, illustrated in Fig. 3(d). “Data Optimization”, the formulation of optimal-data UAP maximin in Eq. 3 is also examined by varying $r$. The results also demonstrate that “Model-Data Optimization” invariably outperforms “Data Optimization”, their optimal $r$ value differ.
5 Limitation and Future Work
Our DM-UAP introduces a novel maximin formulation to enhance the generalization of UAP by dynamically optimizing both the model and data during the UAP generation process. As a result, this incurs additional computational expenses. Specifically, for crafting UAPs on VGG16, the time expense for DM-UAP is approximately 1.6 times that of SGA and twice that of AT-UAP (see Tab. 5). However, considering the universal nature of UAP, these computational costs can be deemed negligible. Once the UAPs are created, there is no need for additional computations, as the off-the-shelf UAPs can be readily utilized to generate AEs at scale. For future studies on crafting UAPs against large models that require more memory, we think adaptively selecting part of the parameters for optimization may be worth exploring.
Method Time consumption Memory consumption Average attack success rate(%)
SPGD 27min’58s 19620 MiB 60.91
AT-UAP 1h’12min’34s 20210 MiB 61.18
SGA 1h’28min’44s 10382 MiB 67.10
Ours 2h’21min’26s 20714 MiB 69.81
Table 5: The optimization expenses of different UAP methods. UAPs are obtained from the VGG16 model.
6 Conclusion
In this paper, we present DM-UAP, a novel approach for generating UAPs with a dynamic maximin optimization strategy. DM-UAP not only optimizes the data used for training but also takes into account dynamic model parameters. Specifically, DM-UAP incorporates an iterative max-min-min optimization framework that dynamically minimizes classification loss to obtain model-data pairs, as well as a curriculum learning algorithm to thoroughly explore the combined landscape of model parameters and data. Extensive experiments on ImageNet demonstrate the superior performance of DM-UAP over state-of-the-art methods, significantly improving the generalization of generated UAPs.
Acknowledgements
Minghui’s work is supported by the National Natural Science Foundation of China (Grant No.62202186). Shengshan’s work is supported by the National Natural Science Foundation of China (Grant No.62372196). Minghui Li is the corresponding author.